# On the Optimization Dynamics of RLVR: Gradient Gap and Step Size Thresholds

### v2

@@ 3.2.1 Convergence and Stagnation
$M(K)\;:\,=\;\sum_{k=0}^{K-1}[\Delta\mu_{q}(k)]_{+}\,\eta_{k},$ (14)
which accumulates the amount of “useful progress” made up to horizon $K$. Intuitively, $M(K)$ grows whenever the update direction is positively aligned with the true objective, and it stagnates when the updates fail to exploit the available signal.
@@
    (Stagnation) Consider when $J_{q}(0)<1$. If the alignment signal is too weak, in the sense that the cumulative alignment remains bounded $M(K)\leq C_{0}$ and $\sum_{k=0}^{\infty}\eta_{k}^{2}\leq C_{0}^{\prime}/(L_{\mathrm{o}}+8G_{\mathrm{o}}^{2})$, for some constants $0\leq C_{0},C_{0}^{\prime}<\infty$, then learning will stall. In this case, the performance remains strictly sub-optimal: $J_{q}(k)\leq J_{q}(0)\big(J_{q}(0)+\exp(C_{0}+C_{0}^{\prime})\,\{1-J_{q}(0)\}\big)^{-1}\!<1$.
@@
    $\eta_{k}\;\leq\;\frac{[\Delta\mu_{q}(k)]_{+}}{2\,(L_{\mathrm{o}}+8\,G_{\mathrm{o}}^{2})}\quad\mbox{where }[\,\cdot\,]_{+}=\max(0,\cdot).$ (15a)
    Then the performance is lower-bounded at any horizon $K$ by
    $J_{q}(K)\;\geq\;\frac{J_{q}(0)}{J_{q}(0)+\{1-J_{q}(0)\}\,\exp\big\{-\frac{1}{2}M(K)\big\}}.$ (15b)
    Moreover, if the alignment accumulates indefinitely, $\lim_{K\rightarrow\infty}M(K)=+\infty$, then the policy is guaranteed to achieve perfect performance: $\lim_{K\rightarrow\infty}J_{q}(K)=1$.
The theorem establishes a clear dichotomy. Convergence is attainable only when update directions exhibit consistent alignment with the underlying objective and the step size is properly scaled to reflect this signal. In the absence of either alignment or adaptive scaling, progress stagnates and the policy remains confined to a suboptimal regime.
@@
which is stated formally in Lemma˜1 of Section˜C.1.1. This inequality shows that $\Delta\mu_{q}(t)\,\eta_{t}$ captures the first-order Taylor approximation of the change in log-odds of $J_{q}$. Summing (16) over iterations and analyzing the resulting terms under different cases reveals that the Cumulative Gap Alignment $M(K)$ governs the value of $J_{q}$. This establishes the claims in Theorem˜1.
3.2.2 The Importance of Properly Chosen Step Size $\eta_{k}$
According to condition (15a) in Theorem˜1(b), the step size $\eta_{k}$ must be carefully scaled to match the gap alignment $\Delta\mu_{q}(k)$. To illustrate this, we contrast two scenarios: a modest step size yields linear convergence, whereas an overly aggressive one causes failure.
@@ 3.2.2 The Importance of Properly Chosen Step Size $\eta_{k}$
Corollary 1 (Linear Convergence with a Uniform Gap).
If every update direction $\boldsymbol{w}_{t}$ provides a uniform gap, $\Delta\mu_{q}(k)\geq\Delta\mu_{q}>0$ for all $k\geq 0$, then a simple fixed step size $\eta$ satisfying
@@
The Perils of Overshooting. The picture changes sharply when the step size is too large. If condition (15a) in Theorem 1(b) is violated, convergence may break down entirely. The next result shows that even with perfect update directions, learning can collapse under overly aggressive step sizes.
@@
There exists a problem instance under ˜1 with $G_{\mathrm{o}}\geq\sqrt{L_{\mathrm{o}}}$ where the Gap Alignment is uniformly positive, $\Delta\mu_{q}(k)\geq\Delta\mu_{q}>0$ for all $k\geq 0$, yet using an overly large constant step size $\eta_{k}=\eta$ leads to failure. Specifically, if the step size satisfies
${\frac{60\,\Delta\mu_{q}}{L_{\mathrm{o}}+G_{\mathrm{o}}^{2}}}\;\leq\;\eta\;\leq\;{\frac{1}{2\sqrt{L_{\mathrm{o}}+G_{\mathrm{o}}^{2}}}}\,$,
@@
Intuition for the lower bound analysis.
Our convergence analysis (Theorem˜1) relies on equation Equation˜16, which uses a first-order approximation of the change in log-odds. For the lower bound, however, it is crucial to examine the second-order expansion. To this end, we define conditional variances over the positive (and negative) response space: ${\rm Var}^{+}:\,={\rm Var}_{\vec{\boldsymbol{o}}\!\,\sim\uppi_{k}(\cdot\mid q,\mathcal{O}_{q}^{+})}\big[\boldsymbol{w}_{k}\cdot\nabla_{\theta}\,\log\uppi_{k}(\vec{\boldsymbol{o}}\!\,\mid q)\big]$. The term $\rm Var$$^{-}$ is defined analogously. The second-order Taylor expansion gives
$\log\Big(\frac{J_{q}(k+1)}{1-J_{q}(k+1)}\Big)-\log\Big(\frac{J_{q}(k)}{1-J_{q}(k)}\Big)\;=\;\Delta\mu_{q}(k)\,\eta_{k}+\{{\rm Var}^{+}-{\rm Var}^{-}\}\cdot\eta_{k}^{2}+\mathcal{O}(\eta_{k}^{3})\,.$ (17)
[…]
@@ 4 Token-Level Analysis
We now move towards a token-level analysis of RLVR, which sharpens the trajectory-level perspective developed earlier. While natural and general for abstract analysis, our analysis in Section˜3 overlooks the autoregressive structure of LLMs: responses are generated token by token, with intermediate Chain-of-Thought (CoT) steps shaping the learning dynamics.
At the trajectory level, the regularity conditions in ˜1 are imposed on the policy score $\nabla_{\theta}\,\log\uppi_{\theta}(\vec{\boldsymbol{o}}\!\,\mid q)$ of the entire response $\vec{\boldsymbol{o}}\!\,$. However, the score can be decomposed into token-wise contributions: $\nabla_{\theta}\,\log\uppi_{\theta}(\vec{\boldsymbol{o}}\!\,\mid q)=\sum_{t=1}^{|\vec{\boldsymbol{o}}\!\,|}\nabla_{\theta}\,\log\uppi_{\theta}(o_{t}\mid q,\vec{\boldsymbol{o}}\!\,_{<t})$, where every token $o_{t}$ requires a forward pass from the language model and thus carries its own regularity properties. This makes it more natural—and ultimately more powerful—to impose assumptions at the token level. Doing so introduces response length as an explicit factor, which will be central to our analysis. Interestingly, as we will see, it also reveals how the training dynamics adapt to task difficulty under the current policy.
We refine ˜1 into the following token-level version.
@@
˜3 characterizes response length: $T_{\infty}$ bounds the worst case, while $T_{\psi_{1}}$ reflects the typical scale. It holds $\mathds{E}_{\uppi_{\theta}}[|\vec{\boldsymbol{o}}\!\,|\mid q]\leq T_{\psi_{1}}\leq T_{\infty}/\log 2$, so that $T_{\psi_{1}}$ may be much smaller than $T_{\infty}$.
With these two assumptions in place, we are ready to present our token-level convergence guarantee. The statement parallels the trajectory-level result, but now incorporates the finer granularity of token-wise dynamics. We retain the key quantities from Section˜3.1, namely the Gap Alignment $\Delta\mu_{q}(k)$ from equation (11), and the Cumulative Gap Alignment $M(K)$ from equation (14).
@@
Complementing the positive result in Theorem˜3, we now show that the step-size scalings with $T_{\infty}$ and $T_{\psi_{1}}$ are essentially tight, as established by the token-level analogue of Theorem˜2 below.
@@
There exists a problem instance satisfying ˜2 with $G_{\mathrm{p}}\geq\sqrt{L_{\mathrm{p}}}$ where the Alignment Gap is always positive, $\Delta\mu_{q}(k)\geq\Delta\mu_{q}>0$, yet choosing a constant step size $\eta_{k}=\eta$ that is too large leads to a complete failure of learning. Specifically, if the step size satisfies
${\frac{120\,\Delta\mu_{q}}{(L_{\mathrm{p}}+G_{\mathrm{p}}^{2})\,T_{\infty}}}\;\leq\;\eta\;\leq\;{\frac{1}{2\sqrt{(L_{\mathrm{p}}+G_{\mathrm{p}}^{2})\,T_{\infty}}}}\,$, (19)
where $0<\Delta\mu_{q}\leq\frac{1}{240}\sqrt{(L_{\mathrm{p}}+G_{\mathrm{p}}^{2})\,T_{\infty}}$, the policy’s performance will strictly decrease at every step, ultimately converging to zero: $J_{q}(k)<J_{q}(k-1)$ and $\lim_{K\rightarrow\infty}J_{q}(K)=0$.
@@
Finally, note that the $(1-J_{q}(k))$ factor only influences how fast convergence proceeds toward 1. In the lower bound construction of Theorem˜4, $J_{q}(k)$ is strictly decreasing, so this term behaves like a constant and does not alter the failure guarantee. Hence, it affects the upper bound but not the lower bound.
@@
We next examine how the update rules of GRPO and Dr. GRPO (or REINFORCE) fit into our token-level framework. For clarity, we restrict attention to the scaling behavior with respect to $\Delta\mu_{q}$, $T_{\psi_{1}}$, $J_{q}$, and $(1-J_{q})$ under a single prompt $q$.
In this regime, the GRPO gradient from equation (5) simplifies to
$\textstyle\boldsymbol{g}_{\mathrm{GRPO}}(\uppi_{k})\;\asymp\;\mathds{E}_{\vec{\boldsymbol{o}}\!\,\sim\uppi_{k}(\cdot\mid q)}\big[A(q,\vec{\boldsymbol{o}}\!\,)\cdot\nabla_{\theta}\,\log\uppi_{k}(\vec{\boldsymbol{o}}\!\,\mid q)\big]\,/\,\big\{T_{\psi_{1}}\sqrt{J_{q}(1-J_{q})}\big\}\,.$
@@
On the other hand, condition (18a) in Theorem˜3, under the simplification $L_{\mathrm{p}}\ll G_{\mathrm{p}}^{2}$ and retaining the $T_{\psi_{1}}/(1-J_{q})$ term in the denominator, reduces to
@@
Sketch of Proof for Theorem˜3.
@@
The formal statement of bound (22) is provided in Lemma˜2 in Section˜C.2. In parallel, we adapt the trajectory-level result (16) to the token setting by taking $G_{\mathrm{o}}=G_{\mathrm{p}}T_{\infty}$ and $L_{\mathrm{o}}=L_{\mathrm{p}}T_{\infty}$. We then combine these two bounds, applying whichever is tighter in a given regime. The remaining steps follow the same structure as in Theorem˜1(b).
@@
5 Numerical Experiments – REINFORCE on Contextual Bandits

### v3

@@ 3.2.1 Convergence and Stagnation
$M_{q}(K)\;:\,=\;\sum_{k=0}^{K-1}[\Delta\mu_{q}(k)]_{+}\,\eta_{k},$ (12)
which accumulates the amount of “useful progress” made up to horizon $K$. Intuitively, $M_{q}(K)$ grows whenever the update direction is positively aligned with the true objective, and it stagnates when the updates fail to exploit the available signal.
@@
    (Stagnation) Consider when $J_{q}(0)<1$. If the alignment signal is too weak, in the sense that the cumulative alignment remains bounded $M_{q}(K)\leq C_{0}$ and $\sum_{k=0}^{\infty}\eta_{k}^{2}\leq C_{0}^{\prime}/(L_{\mathrm{o}}+8G_{\mathrm{o}}^{2})$, for some constants $0\leq C_{0},C_{0}^{\prime}<\infty$, then learning will stall. In this case, the performance remains strictly sub-optimal: $J_{q}(k)\leq J_{q}(0)\big(J_{q}(0)+\exp(C_{0}+C_{0}^{\prime})\,\{1-J_{q}(0)\}\big)^{-1}\!<1$.
@@
    $\eta_{k}\;\leq\;\frac{[\Delta\mu_{q}(k)]_{+}}{2\,(L_{\mathrm{o}}+8\,G_{\mathrm{o}}^{2})}$ (13a)
    where $\cdot\,]_{+}=\max(0,\cdot)$. Then the performance is lower-bounded at any horizon $K$ by
    $\hskip-30.00005ptJ_{q}(K)\;\geq\;\frac{J_{q}(0)}{J_{q}(0)+\{1-J_{q}(0)\}\,\exp\big\{-\frac{M_{q}(K)}{2}\big\}}.$ (13b)
The theorem establishes a clear dichotomy. Convergence is attainable only when update directions exhibit consistent alignment with the underlying objective and the step size is properly scaled to reflect this signal. In the absence of either alignment or adaptive scaling, progress stagnates and the policy remains confined to a suboptimal regime. In the limit, if $\lim_{K\to\infty}M_{q}(K)=+\infty$, then Equation 13b implies we achieve perfect performance $\lim_{K\to\infty}J_{q}(K)=1$. In Appendix G, we show that a converse of this statement is true, showing the divergence of $M_{q}(K)$ tightly characterizes reward convergence.
@@
which is stated formally in Lemma 1 of Subsection C.1.1. This inequality shows that $\Delta\mu_{q}(t)\,\eta_{t}$ captures the first-order Taylor approximation of the change in log-odds of $J_{q}$. Summing (14) over iterations and analyzing the resulting terms under different cases reveals that the Cumulative Gap Alignment $M_{q}(K)$ governs the value of $J_{q}$. This establishes the claims in Theorem 1.
The Importance of Properly Chosen Step Size $\eta_{k}$
According to condition (13a) in Theorem 1(b), the step size $\eta_{k}$ must be carefully scaled to match the Gap Alignment $\Delta\mu_{q}(k)$. To illustrate this, we contrast two scenarios: a modest step size yields linear convergence, whereas an overly aggressive one causes failure.
@@
Corollary 3.1 (Linear Convergence with a Uniform Gap).
If every update direction $\boldsymbol{w}_{k}$ provides a uniform gap, $\Delta\mu_{q}(k)\geq\Delta\mu_{q}>0$ for all $k\geq 0$, then a simple fixed step size $\eta$ satisfying
@@
Uniformly lower bounded Gap Alignments hold in idealized settings (e.g., we provide a toy example of a softmax policy over two responses in Appendix E) but may not hold for real LLMs where the Gradient Gap may vanish as advantages do (Yu et al., 2025). In such situations, our theory (Theorem 6) predicts training stagnation, which has been empirically observed in RLVR.
The Perils of Overshooting. The picture changes sharply when the step size is too large. If condition (13a) in Theorem 1(b) is violated, convergence may break down. The next result shows that even with perfect update directions, learning can collapse under overly aggressive step sizes.
@@
There exists a problem instance under Assumption 1 with $G_{\mathrm{o}}\geq\sqrt{L_{\mathrm{o}}}$ where the Alignment Gap is uniformly positive, $\Delta\mu_{q}(k)\geq\Delta\mu_{q}>0$ for all $k\geq 0$, yet using an overly large constant step size $\eta_{k}=\eta$ leads to failure. Specifically, if the step size satisfies
[…]
@@ 4 Token-Level Analysis
We now derive a token-level analysis of RLVR, which sharpens the trajectory-level perspective developed earlier. While natural for abstract analysis, our analysis in Section 3 overlooks the autoregressive structure of LLMs: responses are generated token by token, with intermediate Chain-of-Thought (CoT) steps shaping the learning dynamics.
At the trajectory level, the regularity conditions in Assumption 1 are imposed on the policy score $\nabla_{\theta}\,\log\uppi_{\theta}(\vec{\boldsymbol{o}}\!\,\mid q)$ of the entire response $\vec{\boldsymbol{o}}\!\,$. However, the score can be decomposed into token-wise contributions: $\nabla_{\theta}\,\log\uppi_{\theta}(\vec{\boldsymbol{o}}\!\,\mid q)=\sum_{t=1}^{|\vec{\boldsymbol{o}}\!\,|}\nabla_{\theta}\,\log\uppi_{\theta}(o_{t}\mid q,\vec{\boldsymbol{o}}\!\,_{<t})$, where every token $o_{t}$ requires a forward pass from the language model and thus carries its own regularity properties. This makes it more natural—and ultimately more powerful—to impose assumptions at the token level. Doing so introduces response length as an explicit factor, which will be central to our analysis. Interestingly, as we will see, it also reveals how the training dynamics adapt to task difficulty under the current policy.
We refine Assumption 1 into a token-level version.
@@
Assumption 3 characterizes response length: $T_{\infty}$ bounds the worst case, while $T_{\psi_{1}}$ reflects the typical scale. It holds $\mathds{E}_{\uppi_{\theta}}[|\vec{\boldsymbol{o}}\!\,|\mid q]\leq T_{\psi_{1}}\leq T_{\infty}/\log 2$, so that $T_{\psi_{1}}$ may be much smaller than $T_{\infty}$.
With these two assumptions in place, we are ready to present our token-level convergence guarantee. The statement parallels the trajectory-level result, but now incorporates the finer granularity of token-wise dynamics. We retain the key quantities from Subsection 3.1, namely the Gap Alignment $\Delta\mu_{q}(k)$ from equation (10), and the Cumulative Gap Alignment $M_{q}(K)$ from equation (12).
@@
Complementing the positive result in Theorem 3, we now show that the step-size scalings with $T_{\infty}$ and $T_{\psi_{1}}$ are essentially tight, as established by the token-level analogue of Theorem 2 below.
@@
There exists a problem instance satisfying Assumption 2 with $G_{\mathrm{p}}\geq\sqrt{L_{\mathrm{p}}}$ where the Gap Alignment is always positive, $\Delta\mu_{q}(k)\geq\Delta\mu_{q}>0$, yet choosing a constant step size $\eta_{k}=\eta$ that is too large leads to a complete failure of learning. Specifically, if the step size satisfies
${\frac{120\,\Delta\mu_{q}}{L_{\mathrm{p}}\cdot T_{\infty}^{1/2}+G_{\mathrm{p}}^{2}\cdot T_{\infty}}}\;\leq\;\eta\;\leq\;{\frac{1}{2\sqrt{L_{\mathrm{p}}\cdot T_{\infty}^{1/2}+G_{\mathrm{p}}^{2}\cdot T_{\infty}}}}\,$, (16)
where $0<\Delta\mu_{q}\leq\frac{1}{240}\sqrt{L_{\mathrm{p}}\cdot T_{\infty}^{1/2}+G_{\mathrm{p}}^{2}\cdot T_{\infty}}$, the policy’s performance will strictly decrease at every step, ultimately converging to zero: $J_{q}(k)<J_{q}(k-1)$ and $\mathop{\lim}_{K\rightarrow\infty}J_{q}(K)=0$.
@@
Finally, note that the $(1-J_{q}(k))$ factor only influences how fast convergence proceeds toward 1. In the lower bound construction of Theorem 4, $J_{q}(k)$ is strictly decreasing, so this term behaves like a constant and does not alter the failure guarantee. Hence, it affects the upper bound but not the lower bound.
@@
We next examine how the update rules of GRPO and Dr. GRPO (or REINFORCE) fit into our token-level framework. In particular, we focus on the scaling behavior with respect to $\Delta\mu_{q}$, $T_{\psi_{1}}$, $J_{q}$, and $(1-J_{q})$ under a single prompt $q$.
In this regime, the GRPO gradient from (5) simplifies to
$\textstyle\boldsymbol{g}_{\mathrm{GRPO}}(\uppi_{k})\;\asymp\;\frac{\mathds{E}_{\vec{\boldsymbol{o}}\!\,\sim\uppi_{k}(\cdot\mid q)}\big[A(q,\vec{\boldsymbol{o}}\!\,)\cdot\nabla_{\theta}\,\log\uppi_{k}(\vec{\boldsymbol{o}}\!\,\mid q)\big]}{T_{\psi_{1}}\sqrt{J_{q}(1-J_{q})}}\,.$
@@
On the other hand, condition (15a) in Theorem 3, under the simplification $L_{\mathrm{p}}\ll G_{\mathrm{p}}^{2}$ and retaining the $T_{\psi_{1}}/(1-J_{q})$ term in the denominator, reduces to
@@
Sketch of Proof for Theorem 3.
@@
The formal statement of bound (19) is provided in Lemma 2 in Subsection C.2. In parallel, we adapt the trajectory-level result (14) to the token setting by taking $G_{\mathrm{o}}=G_{\mathrm{p}}T_{\infty}$ and $L_{\mathrm{o}}=L_{\mathrm{p}}T_{\infty}$. We then combine these two bounds, applying whichever is tighter in a given regime. The remaining steps follow the same structure as in Theorem 1(b).
@@
5 Multi-Prompt Analysis
We next generalize the token-level single-prompt theory developed in Section 4 to the full multi-prompt setup of Section 2. Recall our population-level objective from Equation 2: $J(\uppi_{\theta})\mbox{$:\,=$}\mathds{E}_{q\sim\mathds{P}_{Q},\,\vec{\boldsymbol{o}}\!\,\sim\uppi_{\theta}(\cdot|q)}\big[r^{\star}(q,\vec{\boldsymbol{o}}\!\,)\big]\,.$ We next define batch-analogues of the Gradient Gap and the Gap Alignment.
Joint Positive and Negative Spaces.
We partition the space of prompt-response pairs based on the verifiable reward: $\mathcal{O}^{+}\;:\,=\;\big\{(q,\vec{\boldsymbol{o}}\!\,):r^{\star}(q,\vec{\boldsymbol{o}}\!\,)=1\big\}$ and $\mathcal{O}^{-}\;:\,=\;\big\{(q,\vec{\boldsymbol{o}}\!\,):r^{\star}(q,\vec{\boldsymbol{o}}\!\,)=0\big\}$.
Multi-Prompt Conditional Distributions.
We define the joint distribution over prompt-response pairs under policy $\uppi_{\theta}$ as $P_{\theta}(q,\vec{\boldsymbol{o}}\!\,)\;:\,=\;\mathds{P}_{Q}(q)\cdot\uppi_{\theta}(\vec{\boldsymbol{o}}\!\,\mid q)$. The conditional distributions over positive/negative pairs are defined as:
$\!\!\!P^{+}_{\theta}(q,\vec{\boldsymbol{o}}\!\,)$ $:\,=\frac{\mathds{P}_{Q}(q)\cdot\uppi_{\theta}(\vec{\boldsymbol{o}}\!\,\mid q)}{{J}(\uppi_{\theta})}\cdot\mathds{1}\{(q,\vec{\boldsymbol{o}}\!\,)\in\mathcal{O}^{+}\},$
$\!\!\!P^{-}_{\theta}(q,\vec{\boldsymbol{o}}\!\,)$ $:\,=\frac{\mathds{P}_{Q}(q)\cdot\uppi_{\theta}(\vec{\boldsymbol{o}}\!\,\mid q)}{1-{J}(\uppi_{\theta})}\cdot\mathds{1}\{(q,\vec{\boldsymbol{o}}\!\,)\in\mathcal{O}^{-}\}.$
Multi-Prompt Gradient Gap.
We define the multi-prompt expected score functions over positive and negative pairs ${\boldsymbol{g}}^{+}(\uppi_{\theta}):\,=\mathds{E}_{(q,\vec{\boldsymbol{o}}\!\,)\sim P^{+}_{\theta}}\big[\nabla_{\theta}\,\log\uppi_{\theta}(\vec{\boldsymbol{o}}\!\,|q)\big]$ and ${\boldsymbol{g}}^{-}(\uppi_{\theta}):\,=\mathds{E}_{(q,\vec{\boldsymbol{o}}\!\,)\sim P^{-}_{\theta}}\big[\nabla_{\theta}\,\log\uppi_{\theta}(\vec{\boldsymbol{o}}\!\,|q)\big]$ The batch-level Gradient Gap is defined to be ${\boldsymbol{g}}^{+}(\uppi_{\theta})-{\boldsymbol{g}}^{-}(\uppi_{\theta})$. Consequently, the batch-level Gap Alignment is defined as ${\Delta\mu}(k)\;:\,=\;\boldsymbol{w}_{k}\cdot\big({\boldsymbol{g}}^{+}(k)-{\boldsymbol{g}}^{-}(k)\big)$. where ${\boldsymbol{g}}^{+}(k)$ and ${\boldsymbol{g}}^{-}(k)$ denote the batch-level score functions under $\uppi_{k}$. Define the Cumulative Gap Alignment $M(K)$ as in Equation 12 analogously.
Our main result is then the following convergence guarantee, analogous to Theorem 3.
Theorem 5 (Multi-Prompt Convergence).
Suppose Assumptions 2 and 3 hold uniformly over all prompts in the support of $\mathds{P}_{Q}$, and assume J(0)>0{J}(0)>0. If the step size $\eta_{k}$ satisfies Equation 15a but with the batch-level Gap Alignment ${\Delta\mu}(k)$ replacing $\Delta\mu_{q}(k)$, then:
${J}(K)\;\geq\;\frac{{J}(0)}{{J}(0)+\{1-{J}(0)\}\,\exp\big\{-\frac{1}{2}M(K)\big\}}\,.$ (21)
The extension from single-prompt to batch-level convergence in fact follows almost directly from viewing the batch setting as a single-prompt problem with an empty prompt and augmented response $(q,\vec{\boldsymbol{o}}\!\,)$ which first samples a prompt from the model-free distribution $q\sim\mathds{P}_{Q}$ and then further tokens in the reasoning chain from the policy $\uppi_{\theta}$. This “augmented policy” then has score functions only depending on $\uppi_{\theta}$ as $\mathds{P}_{Q}$ does not depend on $\theta$. Thus, Assumptions 2 and 3, as well as the Gradient Gap and Gap Alignment notions are invariant under this perspective shift.
Implications for GRPO and Dr. GRPO.
We next examine how the update rules of GRPO and Dr. GRPO fit into our batch-level framework. Care is needed when generalizing the single-prompt implications of Theorem 3 to the multi-prompt setting, as GRPO’s per-prompt variance normalization in Equation 5 differs from normalization via the batch variance J(1J){J}(1-{J}). Consequently, the GRPO and Dr. GRPO update directions are not exactly proportional to the batch Gradient Gap vector ${\boldsymbol{g}}^{+}-{\boldsymbol{g}}^{-}$.
Nevertheless, under commonly used dynamic sampling or filtering strategies that exclude prompts with extreme $J_{q}$ values (Yu et al., 2025; Cui et al., 2025), the per-prompt rewards $J_{q}$ remain comparable to the batch rewards J{J} in the sense that $J_{q}/{J}$ is bounded. Under this condition, the per-prompt and batch-level Gradient Gaps are related up to constants, and $\mathds{E}_{q}[\Delta\mu_{q}]\asymp{\Delta\mu}$. This allows us to extend the single-prompt implications to characterize the effective batch-level step sizes of GRPO and Dr. GRPO as in Section 4.
6 Numerical Experiments
6.1 REINFORCE on Contextual Bandits